Weekly LIPAc highlights (19/April-23/April)


The signs of corrosion that occurred in the cooling plate of the Power Supply of the Solenoid 1 were repaired and water circulation restarted. However the cooling plate is degrading fast and procurement of new cooling plates started. It is under assessment when the injector conditioning using 6mm Plasma Electrode can resume.

Radio Frequency (RF) System and RFQ cavity

RF water cooling system was recovered after replacement of the resins in the ion exchange unit. The water quality will be monitored closely during start-up and operation of the RF system. Also for the RFQ skid a rigorous continuous monitoring has been started.

RF injection in the RFQ cavity was resumed mid of this week. After the start-up checks and with the RF-RFQ injection stable around 5% duty cycle, it is planned to let the conditioning running during the night this Friday and then over the weekend.

The repair of the flowmeter of the cooling line of one of two Solid State Power Amplifiers (SSPA) is ongoing (spares have been shipped from the supplier in EU). The MEBT bunchers conditioning will resume after mid-May.

Medium energy Beam Transport (MEBT), MEL (MEBT Extension Line), HEBT (High Energy Beam Transport)                                                                                                                                       

The polarity checks of quadrupole and steering magnetic fields was conducted systematically and successfully completed.

Beam Dump

With strenuous support of CIEMAT team and daily remote sessions the Local Control System integration activities for the BD cooling skid is now almost completed. It was possible to perform first trials to run the BD cooling skid in automatic operation (under local and remote control). The two main pumps of the BD skid are fully operative with pressure and flow autoregulation.

The oxygen control loop was continuously running for about 40 hours to check content reduction, system/control stability and the “long-term” operation (we struggle to reduce oxygen concentration below 90 ppb for reaching at the target value of 1 ppb).

Efforts continue now for debugging controllers, upgrade of the OPI to include all functionalities and completion of MPS-PPS tests.

Control System

Elog server was migrated to DMZ server, dedicated for all services allowing remote participation of EU expert, from LIPAc technical net. Procedure to allow access from internet through SSL-VPN is on-going.

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc.